Check my questions and answers.
1. Does the equation represent a direct variation? If so, find the constant of variation.
5x = -3y
a. no
b. yes; k = -5/3
c. yes; k = 5/3
d. yes; k = -3/5
2. Does the equation represent a direct variation? If so, find the constant of variation.
2x - 4y = 0
a. yes; k = -4
b. no
c. yes; k = 1/2
d. yes; k = -1/2
I choose, D & B. which, I may be wrong.

Answer by Alan3354(69443)

Check my questions and answers.
1. Does the equation represent a direct variation? If so, find the constant of variation.
5x = -3y
a. no
b. yes; k = -5/3
c. yes; k = 5/3
d. yes; k = -3/5
2. Does the equation represent a direct variation? If so, find the constant of variation.
2x - 4y = 0
a. yes; k = -4
b. no
c. yes; k = 1/2
d. yes; k = -1/2
I choose, D & B. which, I may be wrong.
Dun & Bradstreet?
------------------------
For linear equations,
Solve for one variable in terms of the other.
If both are in the numerator, it's direct.
-----------
5x = -3y
x = -3y/5
Direct
--------------------
2x - 4y = 0
2x = 4y
x = 2y
Direct.
===========================
The constant depends on which variable you solve for.
y is usually solved for if it's in x & y
--------
x = -3y/5 --> k = -3/5
y = -5x/3
or y = (-5/3)x --> k = -5/3 (notice it's the reciprocal)
=====================
x = 2y --> k = 2
-----
y = x/2
or y = (1/2)x --> k = 1/2
==============================================================
It's a poorly posed problem (not a rarity).
Should ask for y in terms of x, or vice versa, but didn't.